OpenStudy (anonymous):

How do you find the polynomial function that has the given zeroes? 2, -6

OpenStudy (saifoo.khan):

x = 2, x = -6 (x-2)(x + 6)

OpenStudy (saifoo.khan):

FOIL

OpenStudy (saifoo.khan):

x^2 + 6x - 2x - 12 x^2 + 4x - 12

jimthompson5910 (jim_thompson5910):

x = 2 or x = -6 x-2 = 0 or x +6 = 0 (x-2)(x+6)=0 x^2+4x-12=0 So the function is f(x) = x^2+4x-12